# Contrastive Policy Gradient: Aligning LLMs on sequence-level scores in a supervised-friendly fashion

### v1

@@ Abstract
Reinforcement Learning (RL) has been used to finetune Large Language Models (LLMs) using a reward model trained from preference data, to better align with human judgment. The recently introduced direct alignment methods, which are often simpler, more stable, and computationally lighter, can more directly achieve this. However, these approaches cannot optimize arbitrary rewards, and the preference-based ones are not the only rewards of interest for LLMs (eg., unit tests for code generation or textual entailment for summarization, among others). RL-finetuning is usually done with a variation of policy gradient, which calls for on-policy or near-on-policy samples, requiring costly generations. We introduce Contrastive Policy Gradient, or CoPG, a simple and mathematically principled new RL algorithm that can estimate the optimal policy even from off-policy data. It can be seen as an off-policy policy gradient approach that does not rely on important sampling techniques and highlights the importance of using (the right) state baseline. We show this approach to generalize the direct alignment method IPO (identity preference optimization) and classic policy gradient. We experiment with the proposed CoPG on a toy bandit problem to illustrate its properties, as well as for finetuning LLMs on a summarization task, using a learned reward function considered as ground truth for the purpose of the experiments.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) is a classic finetuning step intended at aligning a Large Language Model (LLM) with human judgment (Ouyang et al., 2022). The underlying principle is to learn a reward model from a preference dataset, and to optimize this reward with a regularized Reinforcement Learning (RL) approach (Fox et al., 2015; Jaques et al., 2017; Geist et al., 2019), usually a Policy Gradient (PG) approach (Williams & Peng, 1991) or a variation like Proximal Policy Optimization (PPO) (Schulman et al., 2017). These methods require on-policy or near on-policy samples, and thus require costly generations from the LLM. They can also be hard to tune and computationally heavy, for example through the use of an additional value network.
@@
We posit that preference-based rewards are not the only rewards worth considering when finetuning an LLM. Not everything can be measured through preferences, which are also costly to label. Such examples are using unit tests as a reward for code generation (Le et al., 2022) or a reward measuring textual-entailment for summarization (Roit et al., 2023). The aim of this paper is to propose an RL approach able to optimize an arbitrary reward while being as simple as direct alignment. It is important to note that we do not introduce any specific reward here, our intent is to provide a convenient and efficient tool for optimizing an arbitrary reward function. In particular, we take inspiration from the contrastive learning objective, which had tremendous success in the self-supervised learning technique (Oord et al., 2018; Chen et al., 2020), and extend it to RL techniques.
@@ 3.1 General objective
This can be seen as a weighted log-likelihood, where the weight is the reward of the generation contrasted with the reward of an independent generation, and its symmetric.
Let $\mu_{1}$ and $\mu_{2}$ be some independent distributions (for example underlying a dataset of pairs of generations), that do not need to be know analytically (contrary to policy gradient with importance sampling), and that can be the same too. The objective to be maximized is then
@@
$L(\pi)$ $=\mathbb{E}_{y\leavevmode\nobreak\ \sim\mu_{1}}[\left(R_{\beta/2}^{\pi}(y)-\overline{R_{\beta/2}^{\pi}}^{\mu_{2}}\right)\ln\frac{\pi(y)}{\pi_{\text{ref}}(y)}]$ (12)
$+\mathbb{E}_{y^{\prime}\leavevmode\nobreak\ \sim\mu_{2}}[\left(R_{\beta/2}^{\pi}(y^{\prime})-\overline{R_{\beta/2}^{\pi}}^{\mu_{1}}\right)\ln\frac{\pi(y^{\prime})}{\pi_{\text{ref}}(y^{\prime})}].$ (13)
@@
$\nabla$ $L(\pi)=\mathbb{E}_{y\leavevmode\nobreak\ \sim\mu_{1}}[\left(R_{\beta}^{\pi}(y)-\overline{R_{\beta}^{\pi}}^{\mu_{2}}\right)\nabla\ln\pi(y)]$ (14)
$+\mathbb{E}_{y^{\prime}\leavevmode\nobreak\ \sim\mu_{2}}[\left(R_{\beta}^{\pi}(y^{\prime})-\overline{R_{\beta}^{\pi}}^{\mu_{1}}\right)\nabla\ln\pi(y^{\prime})].$ (15)
@@ 4 Related works
Contrastive Policy Gradient is related to policy gradient (Williams & Peng, 1991). It can be seen as a sound off-policy policy gradient, but crucially not relying on importance sampling, and thus not requiring clipping techniques such as Proximal Policy Optimization (Schulman et al., 2017), allowing for broader applicability (notably, PPO cannot be applied offline to a dataset of unknown density). This link is even stronger:
@@ 5 Toy experiment
We consider the practical CoPG objective of Eq. (17), for which we recall the gradient Eq. (19). Given that CoPG generalizes policy gradient (PG, Prop. 1), we also experiment with it, to illustrate the importance of choosing the right baseline. For PG, we consider the following sample-based gradient:
@@
For each approach, we train the policy $\hat{\pi}$ with stochastic gradient descent. We use Adam (Kingma & Ba, 2014) with learning rate $10^{-3}$, batches of size 512, and train for 100 epochs. We measure the performance of the trained policy with the regret:
@@
Results are presented in Fig. 1. We can observe that CoPG converges to the right solution, as predicted theoretically (Thm. A). IPO converges to a biased solution. This was to be expected, as it can be shown to optimize the reward $\mathbb{E}_{y^{\prime}\sim\mu_{2}}[\sigma(R(y)-R(y^{\prime})]$ (Azar et al., 2024), which is different from the reward of interest $R$. Regarding policy gradient, we can observe that without baseline, naively applying the policy gradient on off-policy data leads to increase the regret: Learning deteriorates the initial policy. Adding the value baseline helps, but it still converges to a biased solution (and it is an ideal algorithm, as here the baseline is analytically computed, while it has to be estimated in practice). Sample-based CoPG converges to the right solution, showing the importance of choosing the right baseline in an off-policy context.
@@ 6 LLM experiments
Reward Model. Our objective is to provide an approach to optimize arbitrary reward functions. As a proof of concept, for this empirical study, we train a reward model using the preference split of the TL;DR dataset and will consider it as the ground truth reward function to be optimized. We insist right away that we do not claim such a reward model to be the best thing to optimize for improving the LLM, we use it as a proxy for assessing if the proposed approach can indeed optimize a reward at scale. For training the reward function, we use a classic Bradley-Terry model (Bradley & Terry, 1952), optimizing for the loss
@@
Evaluation. Recall that the objective is to know if CoPG can optimize a reward function by learning offline from a fix dataset. To evaluate this, every 50 training steps, we perform generations using the trained model on a fixed batch of 128 prompts from the validation dataset and score them using the reward model. We do the same for IPO and DPO, for which we recall that they are trained for preferences according to the reward model, and not according to the original dataset, for a fair comparison, as the reward model is used for evaluation. We also notice that the reward model was not trained on the validation set, only on the train set.
@@ 6.1 Can CoPG optimize a reward?
For a more complete study, we also provide the related results for IPO and DPO. They do not directly aim at optimizing the reward but the preferences. However, given that, in our case, these preferences are ranked according to the reward model, they should also generate sequences of increasing rewards. For DPO, given that in this specific case the preference follows a Bradley-Terry model, it should indeed optimize the reward function, while IPO optimizes for a different objective, see Azar et al. (2024, Prop. 1 and Thm. 1) for more details.
@@ 6.2 How does CoPG compare to direct alignment?
In this case, the preferences following the Bradely-Terry model (because ranked according to the reward function), DPO should optimize for the reward function, according to Azar et al. (2024, Prop. 1). Yet, it achieves the lowest reward in this experiment. We think that this is due to this convergence result being an asymptotical one. In practice, CoPG uses explicitly the reward signal, while DPO only uses a binarized signal (which completions is preferred), and it would require the algorithm to observe multiple ranking of the same completions to converge to the right solution.
@@ 7 Discussion and perspectives
On a controlled but simple bandit experiment (Sec. 5), we have illustrated empirically the convergence properties of the proposed approach, the importance of choosing the right baseline (the one coming from our derivations rather than the classic value baseline, and even less no baseline), and the advantage of optimizing a reward function rather than preferences derived from these rewards, which leads to a biased solution. On a larger scale LLM experiment (Sec. 6), we have shown that CoPG is able to optimize a reward function, in a fully offline and off-policy manner, conversely to other RL-finetuning approaches, and that it can achieve a higher reward than by using a direct alignment approach on preferences ranked by the reward model.
A core perspective is to study CoPG in an online setting. Indeed, if it works in the offline setting, being the first such RL approach in the context of LLMs (to the best of our knowledge), it is not restricted to this setting. Recent works have highlighted the benefits of using online (or fresh) data for direct alignment (Tang et al., 2024a; Tajwar et al., 2024), and we hypothesize that these findings can also benefit to CoPG. Typically, one can consider using a replay buffer, as classically done in off-policy RL but less common for LLM finetuning. As CoPG uses pairs of generations, we think that this can also open new perspectives by using heterogeneous distributions for the completions, as briefly discussed in Sec. 3.2. For example, one distribution could correspond to the replay buffer, while the other could correspond to exploratory generations (for addressing the exploration/exploitation dilemma), or by using a dataset of good but subotpimal generations, in the spirit of RL from demonstrations. Another important perspective is to experiment CoPG on more tasks and rewards, and we also plan to study its possible extension to the multi-objective RL setting.
@@ Appendix A Proofs of theoretical results
The unique maximizer of $J$ is well known, $J$ being a Legendre-Fenchel transform (eg., see Vieillard et al. (2020, Appx. A)), it is given by
@@
$\beta\ln\frac{\pi(y|x)}{\pi_{\text{ref}}(y|x)}=V(x,y)-\ln Z_{V}(x),$ (33)
with $Z_{V}(x)=\beta\sum_{y}\pi_{\text{ref}}(y|x)\exp\frac{V(x,y)}{\beta}$ the associated (scaled) partition function. In essence, $V(x,y)$ can be understood as the logits of the learnt policy, shifted by the log-probabilites of the reference policy.
@@
=$\stackrel{{\scriptstyle\textit{(a)}}}{{=}}\left(R(x,y)-\frac{1}{2}(V(x,y)-Z_{V}(x))-R(x,y^{\prime})+\frac{1}{2}(V(x,y^{\prime})-Z_{V}(x))\right)\left(V(x,y)-\ln Z_{V}(x)\right)$ (35)
$\;+\left(R(x,y^{\prime})-\frac{1}{2}(V(x,y^{\prime})-Z_{V}(x))-R(x,y)+\frac{1}{2}(V(x,y)-Z_{V}(x))\right)\left(V(x,y^{\prime})-\ln Z_{V}(x)\right)$ (36)

### v2

@@ Abstract
Reinforcement Learning (RL) has been used to finetune Large Language Models (LLMs) using a reward model trained from preference data, to better align with human judgment. The recently introduced direct alignment methods, which are often simpler, more stable, and computationally lighter, can more directly achieve this. However, these approaches cannot optimize arbitrary rewards, and the preference-based ones are not the only rewards of interest for LLMs (e.g., unit tests for code generation or textual entailment for summarization, among others). RL-finetuning is usually done with a variation of policy gradient, which calls for on-policy or near-on-policy samples, requiring costly generations. We introduce Contrastive Policy Gradient, or CoPG, a simple and mathematically principled new RL algorithm that can estimate the optimal policy even from off-policy data. It can be seen as an off-policy policy gradient approach that does not rely on important sampling techniques and highlights the importance of using (the right) state baseline. We show this approach to generalize the direct alignment method IPO (identity preference optimization) and classic policy gradient. We experiment with the proposed CoPG on a toy bandit problem to illustrate its properties, as well as for finetuning LLMs on a summarization task, using a learned reward function considered as ground truth for the purpose of the experiments.
Contrastive Policy Gradient:
Aligning LLMs on sequence-level scores in a supervised-friendly fashion
Yannis Flet-Berliac, Nathan Grinsztajn, Florian Strub, Bill Wu, Eugene Choi, Chris Cremer, Arash Ahmadian, Yash Chandak , Mohammad Gheshlaghi Azar, Olivier Pietquin, Matthieu Geist Cohere
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) is a classic finetuning step intended at aligning a Large Language Model (LLM) with human judgment (Ouyang et al., 2022). The underlying principle is to learn a reward model from a preference dataset, and to optimize this reward with a regularized Reinforcement Learning (RL) approach (Fox et al., 2015; Jaques et al., 2017; Geist et al., 2019), usually a Policy Gradient (PG) approach (Williams and Peng, 1991) or a variation like Proximal Policy Optimization (PPO) (Schulman et al., 2017). These methods require on-policy or near-on-policy samples, and thus require costly generations from the LLM. They can also be hard to tune and computationally heavy, for example through the use of an additional value network.
@@
We posit that preference-based rewards are not the only rewards worth considering when finetuning an LLM. Not everything can be measured through preferences, which are also costly to label. Such examples are using unit tests as a reward for code generation (Le et al., 2022) or a reward measuring textual-entailment for summarization (Roit et al., 2023). The aim of this paper is to propose an RL approach able to optimize an arbitrary reward while being as simple as direct alignment. It is important to note that we do not introduce any specific reward here. Our intent is to provide a convenient and efficient tool for optimizing an arbitrary reward function. In particular, we take inspiration from the contrastive learning objective, which had tremendous success in the self-supervised learning technique (Oord et al., 2018; Chen et al., 2020), and extend it to RL techniques.
@@ 3.1 General objective
This can be seen as a weighted log-likelihood, where the weight is the reward of the generation contrasted with the reward of an independent generation, and it’s symmetric.
Let $\mu_{1}$ and $\mu_{2}$ be some independent distributions (for example underlying a dataset of pairs of generations), that do not need to be known analytically (contrary to policy gradient with importance sampling), and that can be the same too. The objective to be maximized is then
@@
$L(\pi)$ $=\mathbb{E}_{y~{}\sim\mu_{1}}[\left(R_{\beta/2}^{\pi}(y)-\overline{R_{\beta/2}^{\pi}}^{\mu_{2}}\right)\ln\frac{\pi(y)}{\pi_{\text{ref}}(y)}]$ (12)
$+\mathbb{E}_{y^{\prime}~{}\sim\mu_{2}}[\left(R_{\beta/2}^{\pi}(y^{\prime})-\overline{R_{\beta/2}^{\pi}}^{\mu_{1}}\right)\ln\frac{\pi(y^{\prime})}{\pi_{\text{ref}}(y^{\prime})}].$ (13)
@@
$\nabla$ $L(\pi)=\mathbb{E}_{y~{}\sim\mu_{1}}[\left(R_{\beta}^{\pi}(y)-\overline{R_{\beta}^{\pi}}^{\mu_{2}}\right)\nabla\ln\pi(y)]$ (14)
$+\mathbb{E}_{y^{\prime}~{}\sim\mu_{2}}[\left(R_{\beta}^{\pi}(y^{\prime})-\overline{R_{\beta}^{\pi}}^{\mu_{1}}\right)\nabla\ln\pi(y^{\prime})].$ (15)
@@ 3.2 A simple sample-based objective
To further simplify the practical implementation, one can easily verify that
$\nabla\ell_{\texttt{CoPG}}(y,y^{\prime};\pi)\propto-\nabla\left(R_{\beta}^{\pi}(y)-R_{\beta}^{\pi}(y^{\prime})\right)^{2},$ (20)
We provide the related pseudocode (used in our experiments) in Alg. 1.
input : Dataset $\mathcal{D}=\{(x,y,y^{\prime})\}$, reference model $\pi_{\text{ref}}$, model to train $\pi_{\theta}$
Initialize $\pi_{\theta}\leftarrow\pi_{\text{ref}}$;
for $t\leftarrow 1$ to $T$ do
       Sample a batch $\mathcal{B}$ from $\mathcal{D}$;
       Define the regularized reward;
       $R_{\beta}^{\pi}(x,y)=R(x,y)-\beta\ln\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$;
       Compute the gradient;
       $g_{t}=\frac{1}{|\mathcal{B}|}\sum_{x,y,y^{\prime}\in\mathcal{B}}\nabla\left(R_{\beta}^{\pi}(x,y)-R_{\beta}^{\pi}(x,y^{\prime})\right)^{2}$;
       Update the parameters;
       $\theta_{t+1}\leftarrow\theta_{t}-\eta_{t}\nabla g_{t}$;
Algorithm 1 Practical offline CoPG
@@ 4 Related works
Contrastive Policy Gradient is related to policy gradient (Williams and Peng, 1991). It can be seen as a sound off-policy policy gradient, but crucially not relying on importance sampling, and thus not requiring clipping techniques such as Proximal Policy Optimization (Schulman et al., 2017) or variations (Wu et al., 2023), allowing for broader applicability (notably, PPO cannot be applied offline to a dataset of unknown density). This link is even stronger:
@@ 5 Toy experiment
We consider the practical CoPG objective of Eq. (17), for which we recall the gradient Eq. (19). Given that CoPG generalizes policy gradient (PG, Prop. 1), we also experiment with it to illustrate the importance of choosing the right baseline. For PG, we consider the following sample-based gradient:
@@
For each approach, we train the policy $\hat{\pi}$ with stochastic gradient descent. We use Adam (Kingma and Ba, 2014) with learning rate $10^{-3}$, batches of size 512, and train for 100 epochs. We measure the performance of the trained policy with the regret:
@@
Results are presented in Fig. 1. We can observe that CoPG converges to the right solution, as predicted theoretically (Thm. A). IPO converges to a biased solution. This was to be expected, as it can be shown to optimize the reward $\mathbb{E}_{y^{\prime}\sim\mu_{2}}[\sigma(R(y)-R(y^{\prime})]$ (Azar et al., 2024), which is different from the reward of interest $R$. Regarding policy gradient, we can observe that without baseline, naively applying the policy gradient on off-policy data leads to an increase in the regret: Learning deteriorates the initial policy. Adding the value baseline helps, but it still converges to a biased solution (and it is an ideal algorithm, as here the baseline is analytically computed, while it has to be estimated in practice). Sample-based CoPG converges to the right solution, showing the importance of choosing the right baseline in an off-policy context.
@@ 6 LLM experiments
Reward Model. Our objective is to provide an approach to optimize arbitrary reward functions. As a proof of concept, for this empirical study, we train a reward model using the preference split of the TL;DR dataset and will consider it as the ground truth reward function to be optimized. We insist right away that we do not claim such a reward model to be the best thing to optimize for improving the LLM, we use it as a proxy for assessing if the proposed approach can indeed optimize a reward at scale. For training the reward function, we use a classic Bradley-Terry model (Bradley and Terry, 1952), optimizing for the loss
@@
Evaluation. Recall that the objective is to know if CoPG can optimize a reward function by learning offline from a fixed dataset. To evaluate this, every 50 training steps, we perform generations using the trained model on a fixed batch of 128 prompts from the validation dataset and score them using the reward model. We do the same for IPO and DPO, for which we recall that they are trained for preferences according to the reward model, and not according to the original dataset, for a fair comparison, as the reward model is used for evaluation. We also note that the reward model was not trained on the validation set, only on the train set.
@@ 6.1 Can CoPG optimize a reward?
For a more complete study, we also provide the related results for IPO and DPO. They do not directly aim at optimizing the reward but rather at optimizing the preferences. However, given that, in our case, these preferences are ranked according to the reward model, they should also generate sequences of increasing rewards. For DPO, given that in this specific case the preference follows a Bradley-Terry model, it should indeed optimize the reward function, while IPO optimizes for a different objective, see Azar et al. (2024, Prop. 1 and Thm. 1) for more details.
@@ 6.2 How does CoPG compare to direct alignment?
In this case, the preferences following the Bradely-Terry model (because ranked according to the reward function), DPO should optimize for the reward function, according to Azar et al. (2024, Prop. 1). Yet, it achieves the lowest reward in this experiment. We think that this is due to this convergence result being an asymptotical one. In practice, CoPG uses explicitly the reward signal, while DPO only uses a binarized signal (which completions is preferred), and it would require the algorithm to observe multiple rankings of the same completions to converge to the right solution.
@@
6.3 KL-Reward trade-off
In regularized RL, the objective is to maximize the reward while keeping the KL divergence between the policy and the reference model small enough. This is especially important for LLMs, as this can prevent reward hacking for example. For this experiment, we consider the HH dataset from Bai et al. (2022). We use the same recipe as TL;DR for training the SFT checkpoint (this time using the preferred examples, in the absence of SFT data), the reward model, IPO, DPO and CoPG. We also use the same hyperparameters.
Figure 7: KL-Reward trade-off for CoPG, DPO and IPO.
We follow the same methodology, sweeping for each algorithm over values of $\beta$, selecting the one with the higher final reward averaged of 128 generations. Then, we retrain each algorithm with its best value of $\beta$, using 1024 generations for computing the reward at each validation step. From that, we compute the KL/reward Pareto front for IPO, DPO and CoPG.
Results are presented in Fig. 7. It is a scatter plot of the rewards as a function of the KL divergences, each point corresponding to one of the 1024 generations for a specific validation step and a specific algorithm. The figure also provides a polynomial interpolation. The Pareto front clearly shows that CoPGoutperforms IPO and DPO in achieving higher rewards while maintaining lower KL divergence. In particular, we note that DPO tends to achieve a higher KL more easily than other methods.
@@ 7 Discussion and perspectives
On a controlled but simple bandit experiment (Sec. 5), we have illustrated empirically the convergence properties of the proposed approach, the importance of choosing the right baseline (the one coming from our derivations rather than the classic value baseline, and even less no baseline), and the advantage of optimizing a reward function rather than preferences derived from these rewards, which leads to a biased solution. On a larger scale LLM experiment (Sec. 6), we have shown that CoPG is able to optimize a reward function, in a fully offline and off-policy manner, conversely to other RL-finetuning approaches, and that it can achieve a higher reward than by using a direct alignment approach on preferences ranked by the reward model. Our experimental study also suggests that this higher reward is achieved at a lower KL cost, compared to IPO or DPO.
A core perspective is to study CoPG in an online setting. Indeed, if it works in the offline setting, being the first such RL approach in the context of LLMs (to the best of our knowledge), it is not restricted to this setting. Recent works have highlighted the benefits of using online (or fresh) data for direct alignment (Tang et al., 2024a; Tajwar et al., 2024), and we hypothesize that these findings can also benefit to CoPG. Typically, one can consider using a replay buffer, as classically done in off-policy RL but less common for LLM finetuning. As CoPG uses pairs of generations, we think that this can also open new perspectives by using heterogeneous distributions for the completions, as briefly discussed in Sec. 3.2. For example, one distribution could correspond to the replay buffer, while the other could correspond to exploratory generations (for addressing the exploration/exploitation dilemma), or by using a dataset of good but suboptimal generations, in the spirit of RL from demonstrations. Another important perspective is to experiment CoPG on more tasks and rewards, and we also plan to study its possible extension to the multi-objective RL setting.
@@ Appendix A Proofs of theoretical results
The unique maximizer of $J$ is well known, $J$ being a Legendre-Fenchel transform (e.g., see Vieillard et al. (2020, Appx. A)), it is given by
@@
$\beta\ln\frac{\pi(y|x)}{\pi_{\text{ref}}(y|x)}=V(x,y)-Z_{V}(x),$ (34)
with $Z_{V}(x)=\beta\ln\sum_{y}\pi_{\text{ref}}(y|x)\exp\frac{V(x,y)}{\beta}$ the associated (scaled) partition function. In essence, $V(x,y)$ can be understood as the logits of the learnt policy, shifted by the log-probabilites of the reference policy.
@@
=$\stackrel{{\scriptstyle\textit{(a)}}}{{=}}\left(R(x,y)-\frac{1}{2}(V(x,y)-Z_{V}(x))-R(x,y^{\prime})+\frac{1}{2}(V(x,y^{\prime})-Z_{V}(x))\right)\left(V(x,y)-Z_{V}(x)\right)$ (36)
$\;+\left(R(x,y^{\prime})-\frac{1}{2}(V(x,y^{\prime})-Z_{V}(x))-R(x,y)+\frac{1}{2}(V(x,y)-Z_{V}(x))\right)\left(V(x,y^{\prime})-Z_{V}(x)\right)$ (37)
@@
Appendix B Qualitative results
We list two examples from the TL;DR dataset during evaluation. We observe that the summary improves in quality during training, with more elaborate structures. Additionally, we started noticing some reward hacking artifacts after 2 epochs ($2\times 600$ steps) of the training.
SUBREDDIT: relationships TITLE: (24/F) I hate losing arguments to my bf (27/m) POST: I just got into another heated petty argument with my bf of 3 years! He started telling our roommate a story about someone, I was sure he was thinking of something else, he stops dead in his tracks and tells me that that’s not what happened. I said more details of what I remembered, he told me again that I was wrong and I should believe him because it’s a story that happened to him. He said what he always does, "why is it that when we disagree you assume I must be wrong? Why don’t you ever consider that you could be wrong?" By that point I had realized it was possible I was wrong, but I absolutely did not want to concede that to him or offer him any slack. I don’t get like this with everybody, but my bf and I both have this stupid trait of never backing down! And I hate admitting I’m wrong to him because I feel like he acts super smug and condescending, when I make a mistake he laughs at me in an "oh, you" sort of way that INFURIATES me. I know I overreact to this sort of thing but I don’t know how to improve!
DPO:
  • Step=300: I hate admitting to my bf that I’m wrong and I get super defensive when he acts like he’s so happy and smug that I was wrong. How do I improve?
  • Step=600: I hate admitting I’m wrong to my bf because he acts super smug and I don’t like it. How do I get better about this?
  • Step=900: I hate admitting I’m wrong to my bf in arguments because he acts super smug and I can’t stand it, but I also hate losing arguments so I always fight to the death with him even when I’m sure I’m wrong sometimes! What do I do?!
  • Step=1200: I hate admitting I’m wrong to my bf because he acts super smug and I want to win every argument. How do I improve ?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!!?!?!?!?!!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!?!
IPO:
  • Step=300: I hate admitting to my bf that I’m wrong and I get super defensive when he acts like he’s so happy that he was right and I was wrong. I want to improve so I don’t get so upset over this!
  • Step=600: I hate admitting I’m wrong to my bf because he acts super smug and I don’t like it, but I also hate losing arguments and always want to be right. I need to learn how to not care so much!
  • Step=900: I hate admitting I’m wrong to my bf in arguments because he acts super smug and I can’t stand it, but I also hate losing arguments! How do I improve so I don’t feel so defensive about being wrong?
  • Step=1200: I hate admitting I’m wrong to my bf in arguments because he acts super smug and I want to win every argument anyway even if I’m not 100% sure I’m right. How do I stop?
CoPG:
  • Step=300: I hate admitting I’m wrong to my bf in arguments, he always acts super smug and I can’t stand it. How can I get better about this?
  • Step=600: I hate admitting I’m wrong to my bf and I hate when he acts super smug and condescending when he knows he has proven he was right and I was wrong. I need to learn how to not let this get to me!
  • Step=900: I hate admitting defeat in petty arguments to my bf, he always acts so smug and I can’t stand it, but I don’t know how to improve so I don’t get as upset with him over it!
  • Step=1200: I hate admitting that I’m wrong to my bf after a disagreement because he acts like a superior being, I don’t want to back down and admit it to him, but I also don’t want to continue the argument because it just gets more heated and nasty.
SUBREDDIT: relationships TITLE: Me [19 F] with my boyfriend [19 M ] 3.5 years, when is it time to let go? with line breaks. POST: Okay, so I’ve never really posted anything before, so sorry if this is weird or I’m not doing it right or something. I started dating my current BF when I was 15 and a sophomore in high school, he was 16 and a junior. Now we are both in college. We are insanely close. Like spend literally ALL our time together. He is incredibly important to me and we do everything together. I love him a lot and I have incredible respect for him as a person. The problem is we both don’t have friends outside of the relationship. I mean we both have friends. We have a lot of friends really, but we just don’t feel like we connect with anyone other than each other anymore. It’s hard to explain. We both ditched all of our friends from high school. Some of it was because of our relationship, but a lot of it was high school is shitty.
DPO:
  • Step=300: I know I need to break up with my boyfriend of 3.5 years, but I’m scared because I feel like I won’t ever connect with anyone again because we have kind of abandoned our friends for each other.
  • Step=600: I know I need to break up with my boyfriend of 3.5 years, but I feel like I won’t ever connect with anyone else because I’ve been so focused on this relationship my whole life. I don’t know what to do.
  • Step=900: I know I need to break up with my boyfriend I’ve been with for 3.5 years, but I’m afraid because I feel like I won’t ever connect with anyone again after him. I feel like I only connect with him. I don’t feel like I connect with anyone else.
  • Step=1200: My boyfriend and I have pretty much zero outside friends and we are too dependent on each other, I know we should break up but I’m scared I will never connect with anyone else. I’d really appreciate some advice. Thanks.
IPO:
  • Step=300: I know I need to break up with my boyfriend of 3.5 years, but I’m scared because I feel like I won’t ever connect with anyone outside of my SO now. Has anyone expressed something like this?
  • Step=600: I know I need to end my relationship with my boyfriend whom I love very dearly, but I’m scared because I feel like I won’t ever connect with anyone outside of the relationship anymore if I do.
  • Step=900: I know it’s time to move on from my long-term relationship, but I’m scared because I don’t feel like I’ll ever connect with anyone again outside of my SO. Has anyone else expressed this feeling before? If so, what did you do?
  • Step=1200: My boyfriend and I have pretty much ditched all of our friends from high school and replaced them with each other. We know we need to break up eventually, but I’m scared because I don’t feel like I’ll ever connect with anyone else. Has anyone else experienced something like this?
CoPG:
  • Step=300: I know I need to break up with my boyfriend of 3.5 years whom I love very dearly, but I’m scared because I feel like I won’t ever connect with anyone else outside of this relationship. Has anyone expressed this feeling before?
  • Step=600: I know I need to end my relationship with my boyfriend whom I love very dearly, but I am scared I’ll never find anyone I connect with after him because I feel like I don’t connect with anyone other than him.
  • Step=900: I know I need to end my long-term relationship, but I’m scared because I don’t feel like I’ll ever be able to connect with anyone else because all of my friends currently are either from this relationship or I made this year and don’t feel like I connect with them.
  • Step=1200: My boyfriend and I are pretty much all each other have ever cared about or focused on in terms of friends. I know I need to break up with him, but I’m scared I won’t find any friends I connect with ever again.